-9+8(9x-8)=-1+7(12x+12)
1
Expert's answer
2012-10-26T11:31:11-0400
-9+8(9x-8)=-1+7(12x+12)
-9 + 8*9x – 8*8 = -1 + 7*12x + 7*12
72x – 73 = 84x + 83
72x – 84x = 83 + 73
-12x = 156
X = -13